Simple random sampling uses a sample of size n from a population of size N to obtain data that can be used to make inferences about the characteristics of a population. Suppose that, from a population of 60 bank accounts, we want to take a random sample of nine accounts in order to learn about the population. How many different random samples of nine accounts are possible?
Answer #1

Answer)

We can estimate the answer by the formula of combinations

If we want to select r things from a total of n things

Then total possible outcomes is given by ncr = n!/(r!*(n-r)!)

N! = n*n-1*n-2*n-3*n-4.....till 1

For example 5! = 5*4*3*2*1

Here n = 60

R = 9

60c9 = 60!/(9!*(60-9)!) = 14783142660